Homework: 2-2 MyStatLab: Module Two Problem Set 12 of 15 (11 D Any sample sine oouid be uned
Answer #1

Solution(a)
Its answer is B. i.e sample size required greater than or equal to 30.

Solution(b)
Given n = 35
P(Xbar<15)
Z = (16.3-15)/4.8/sqrt(35)
Z = 1.3/4.8/5.92 = 1.60
From Z table we found p-value
P(Xbar<15) = 0.9452
So there is 94.52% probability that is sample mean time is less than 15 minutes.